OpenStudy (anonymous):

3 - 5 + 3b - 13 ???? show work please

OpenStudy (anonymous):

pemdas: 3-5+3b-13 combine like terms 3b=-3+5+13 moved all to one side with variable staying on its side. 3b= 2+13 3b=15 divide both sides by 3 (3b)/3=15/3 b=5